KaydenKroupa:

The manufacturing cost of a cell phone was $136. Sellers first marked up the cost by 25 percent. Sales were poor, so sellers decreased the markup to 17 percent. Calculate the old and new selling prices. selling price = original cost + markup value A. old price was $170; new price is $164.90 B. old price was $150.12; new price is $117 C. old price was $170; new price is $159.12

sukie:

Do you know how to get the percentage cost?

sukie:

100% =====> $136 25% =====> ?

sukie:

I used the cross method so it's like this : Old price = manufacturing cost + (percentage*manufacturing cost)/100 Old price =136+(25*136)/100 Old price =$170

sukie:

the other part is simple same steps. just change the percentage 25 to 17 in the equation above.

sukie:

New price =136+(17*136)/100 New price =$159.12